• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name saha

    Log entry time 08:03:18 on November 3,1998

    Entry number 4587

    keyword=Arc energy measurement

    We have just finished an energy-arc measurement. We had several problems, the most serious of which is that the ARC PMT high voltages were zero and Yves had to perform some black magic to set them to the right values. This needs to be documented better and he is writing a followup halog message regarding this.

    At 0200, we asked MCC for setting up dispersive mode to do the Arc energy measurement. MCC was finally ready at 0630 delivering the right dispersive optics without any trips to our hall. As we did a harp scan, the 3rd harp (1HA1C18A) tripped the Compton ion chamber. Since we were not going through the Compton chicane, I gave permission for them to mask only the Compton ion chamber. At this point we were also noticing absence of any peaks in our harp scans which then led us to our main problem i.e. that the ARC PMT high voltages were all zero. After Yves fixed this problem with some kluges, we got some good harp scans with 5 microA cw. The gains had to be adjusted and the best scans are in files: nov3-arc-5 and nov3-arc-6 with pmt gains of 4/4/16/32. A picture is appended below. We also did a dipole field measurement, the file is in integral_250.data. The quality of the data seems good, and Pascal needs to analyse the data now.


    FIGURE 1